OpenStudy (anonymous):

can some one help me please? pretty please ? 4,8,5,12,3,9,5,2 find the nearest hundredth and the mean and standard deviation please please ?

OpenStudy (calculusfunctions):

Assuming population data, the mean\[\mu = \frac{ \sum_{}^{}x _{i} }{ N }\]where \[x _{1}, x _{2}, x _{3}, x _{4}, ... x _{i}\]is the given data and N is the total number of data. Standard deviation\[\sigma = \sqrt{\frac{ \sum_{}^{}(x -\mu)^{2} }{ N }}\] For sample data, replace N with n - 1, where n - 1 is one less than the total number of data. Also for sample data, we use x bar as the symbol for mean instead of the Greek letter mu and we use s as the symbol for standard deviation instead of the Greek letter sigma (lower case).